mathslover (mathslover):

Application of Derivatives... Problem... Prove that ln (1+x) < x for x > 0 .

mathslover (mathslover):

@BSwan and @ganeshie8

OpenStudy (anonymous):

so u need to prove that ln (1+x) < x x - ln (1+x)>=0 or ln (1+x)-x<0

mathslover (mathslover):

Yep, I think we need to prove \(\ln(1+x) - x < 0\) or \(-\ln (1+x) + x > 0\)

mathslover (mathslover):

Okay, got it.

OpenStudy (anonymous):

yep , ok how did u got it ?

mathslover (mathslover):

f(X) = ln(1+x) - x f'(x) = \(\cfrac{1}{1+x} - 1\) \(\cfrac{1 - (1+x)}{1+x}\) \(\cfrac{-x}{1+x}\) now as f'(x) > 0 (for x >0) therefore, \(\cfrac{-x}{1+x} < 0\) (for x > 0)

mathslover (mathslover):

f(x) is thus, decreasing for x > 0. f(x) < f(0) \(\ln(1+x) - x < f(0)\) \(\ln (1+x) < x\)

mathslover (mathslover):

This is how I did it...

OpenStudy (anonymous):

well mmm but it not tells u it dnt cross the x axis :O 

mathslover (mathslover):

Why does it need to cross the x-axis? (sorry, I am yet to learn the geometrical representations in this chapter)

OpenStudy (anonymous):

well if its >0 then it dnt cross the x axis

mathslover (mathslover):

Okay yeah as x > 0 then it should not cross the x axis...

mathslover (mathslover):

Is that graph for ln(1+x) < x (for all x > 0) ?

OpenStudy (anonymous):

@mathslover, your method is right. I would have done it slightly differently. Let \(f(x)=\ln(1+x)\) and \(g(x)=x\). We know that \(f(0)=g(0)=0\), so to show that \(f(x)<g(x)\) we need to show that \(g(x)\) is increasing faster than \(f(x)\). To do so, we must show \(f'(x)<g'(x)\): \[\begin{cases}f'(x)=\dfrac{1}{1+x}\\\\g'(x)=1\end{cases}\] Clearly, for all \(x>0\), we have \(\dfrac{1}{1+x}<1\) since \(\dfrac{1}{1+x}\to0\) as \(x\to\infty\). And we're done.

mathslover (mathslover):

@SithsAndGiggles -> That was a great method. I appreciate that! Thanks a lot.